### v2

# Anomalous Energy Injection in the Gross-Pitaevskii Framework for Turbulence in Neutron Star Glitches

Abstract
Neutron star glitches—sudden increases in rotational frequency—are thought to result from angular momentum transfer via quantized vortices in the superfluid core. To investigate the underlying superfluid dynamics, we employ a two-dimensional rotating atomic Bose-Einstein condensate described by a damped Gross-Pitaevskii equation with an imposed pinning potential that serves as a simplified analogue of a crust. Within this minimal framework, we examine the emergence and evolution of turbulent vortex motion following impulsive perturbations reminiscent of glitch-like forcing. Our simulations reveal a transient Kolmogorov-like turbulent cascade ($k^{-5/3}$) that transitions to a Vinen-like scaling ($k^{-1}$). We identify an anomalous secondary injection mechanism driven primarily by quantum pressure, which can sustain turbulent fluctuations in such a system. By tuning the damping coefficient $\gamma$, we determine an optimal regime for energy transfer. While idealized, these findings illustrate how quantum turbulence with multiple scaling regimes can arise in pinned, rotating superfluids, and they suggest possible qualitative connections to vortex-mediated dynamics in neutron stars and other astrophysical superfluid systems.
Introduction: Decaying neutron stars exhibit sudden increases in rotation frequency, known as glitches [1], caused by the transfer of angular momentum from quantized vortices in the neutron superfluid to the outer crust of the pulsar [2, 3]. This transfer is primarily triggered by the vortex avalanche mechanism, in which quantum vortices become pinned to crustal nuclei [4], preventing the superfluid from spinning down at the same rate as the crust. The resulting rotational lag between the crust and superfluid generates a Magnus force on the vortices [5]. Once a critical velocity is reached, the quantized vortices unpin from their pinning sites [6, 7]. These vortices are then expelled toward the crust, transferring their angular momentum and producing a glitch [8]. This glitch process involves a complex vortex avalanche encompassing millions of vortices, triggering their depinning [9].
Modeling the depinning avalanche mechanism poses significant challenges. Vortex depinning initiates when a vortex, initially trapped by a pinning site, overcomes the attractive force and begins moving in the direction of the applied superfluid flow. Several factors, such as the critical velocity of the superfluid flow and the size and shape of the pinning site, affect the overall dynamics of depinning [10]. However, an alternative analysis suggests that glitches emerge due to a quantum transition in which the core transforms into a superconducting gluon-quark superfluid state, with resultant vortex depinning occurring at the boundary layer between the core and the surrounding dissipative medium [3]. In this Letter, we demonstrate anomalous energy transfer from quantum pressure to incompressible kinetic energy, thereby proposing a candidate mechanism for turbulent dynamics within a Gross-Pitaevskii analog of neutron star glitches.
Rotating Bose-Einstein condensates (BECs) provide a quantitatively precise, experimentally tunable, and microscopically resolvable analogue for investigating vortex avalanche processes in neutron-star superfluids [11, 12, 13]. Verma et al. [14] and Shukla et al. [15] developed a minimal model for the emergence of the superfluid glitches, based on the interaction of the neutron-superfluid vortices and proton-superconductor flux tubes. More recently, Poli et al. [16] extended Gross–Pitaevskii models of pinned superfluids by including dipolar interactions and realistic crustal pinning potentials, demonstrating the emergence of self-organised criticality in simulated glitch events. Although these and related Gross–Pitaevskii–based approaches have successfully reproduced many observed features of pulsar glitches, they remain limited by the modest number of vortices and the small physical size of the simulated condensates; limitations that are unavoidable with current computational resources and that do not affect more phenomenological (non-microscopic) vortex-avalanche models. Building on this Gross–Pitaevskii framework, our study investigates the turbulent superfluid dynamics triggered by large vortex avalanches, uncovering a previously overlooked anomalous energy-injection mechanism.
The inherent compressibility of BECs renders them susceptible to quantum turbulence (QT) under dynamic instabilities. This turbulence is primarily characterized by the breakdown of vortices in a self-similar process, leading to a power-law scaling of $k^{-5/3}$, known as the Kolmogorov spectrum [17, 18]. In addition to the Kolmogorov regime, QT in BECs also manifests in the Vinen regime, which is characterized by random vortex distributions in superfluids and exhibits a $k^{-1}$ scaling. Recent developments have identified novel turbulence regimes, including strong quantum turbulence [19, 20], marked by density fluctuations and non-polarized vortex lines, as well as rotational quantum turbulence [21, 22, 23, 24], and turbulence in self-gravitating dark-matter BEC candidates [25, 26].
Previous studies have explored the nature of glitches and vortex avalanches in the context of pinning potentials and superfluid behavior. However, the role of superfluid-crust interactions in inducing turbulent flow remains largely unexamined. In this letter, we investigate the characteristics of turbulence and associated velocity flows arising from such interactions. Furthermore, it analyzes the mechanisms driving this turbulence, facilitated by pinning sites and the dynamic spin-down of the condensate. The damping effects on the system are also evaluated, identifying an optimal damping coefficient that maximizes turbulence strength. Recent studies indicate that quantum turbulence suppresses collective excitation modes in BECs [27, 28]. As these modes drive vortex avalanches and glitch phenomena, quantifying the onset of turbulence in such systems is essential. Beyond its implication for existing neutron star models, the effect of vortex depinning, particularly its role in driving incompressible turbulence through secondary injection, is of significant interest in the domain of quantum turbulence [29, 30, 31, 32, 19].
The Model: We model the interior of a spinning-down neutron star’s superfluid core using a quasi-two-dimensional (quasi-2D) atomic Bose-Einstein condensate. This rotating BEC system, incorporating dissipative interactions, is described by the damped Gross-Pitaevskii equation (dGPE), presented in its non-dimensional form in Ref. [33].
$(\mathrm{i}-\gamma)\frac{\partial\psi}{\partial t}=\left[-\frac{1}{2}\nabla^{2}+V(\mathbf{r},t)+g|\psi|^{2}-\Omega(t)L_{z}\right]\psi,$ (1)
where $\psi\equiv\psi(\mathbf{r},t)$ is the condensate wave function, with $\mathbf{r}\equiv(x,y)$. $\nabla^{2}=\partial_{x}^{2}\,+\partial_{y}^{2}$ represents the two-dimensional Laplacian, $\gamma$ corresponds to the parameterized damping coefficient, and $g$ is the nonlinear interaction strength given by $g=800$. Although the interaction parameter does not directly represent the microscopic nuclear interactions in a neutron superfluid, it models a short-range repulsive potential that captures the essential physics for generating quantum vortices. For the large-scale hydrodynamic behavior and collective vortex dynamics relevant to glitch-scale turbulence, the system can be described by effective theories where microscopic details are parameterized (e.g., into a mutual friction coefficient), rather than acting as the primary driver of the dynamics [34].
The time-dependent rotation frequency $\Omega(t)$ describes the rotational deceleration profile of the condensate as follows:
$\Omega(t)=\begin{cases}\Omega_{0}\cos^{2}\left(\frac{\pi t}{2t_{s}}\right),&\text{if }t\leq t_{s},\\ 0,&\text{if }t>t_{s},\end{cases}$ (2)
where $\Omega_{0}$ is the initial rotation frequency of the condensate, and $t_{s}$ is the spin-down time. The $z$-component of the angular momentum is given by $L_{z}=\mathrm{i}\hbar(y\partial_{x}-x\partial_{y})$. In general, glitch events arise from transitions between two distinct non-zero rotational frequencies [35, 16]. However, a partial spin-down can retain vortices (see supplementary Fig.2), which may suppress the amplitude of glitches. While introducing a sudden deceleration can overcome this suppression, it often results in an unstable and unconfined condensate. Therefore, to enhance the vortex avalanche process without destabilizing the condensate, we have instead chosen to fully spin down the condensate.
The potential term $V(\mathbf{r},t)$ comprises the following components
$V(\mathbf{r},t)=V_{\mathrm{box}}(\mathbf{r})+V_{\mathrm{crust}}(\mathbf{r},t)+V_{\mathrm{cent}}(\mathbf{r},t).$ (3)
The confining circular-box potential, $V_{\mathrm{box}}$, with radius $R_{\mathrm{box}}$, is defined as $V_{\mathrm{box}}(\mathbf{r})=V_{0b}\Theta(r-R_{\mathrm{box}})$ [33], where $\Theta$ is the Heaviside step function and $r=\sqrt{x^{2}+y^{2}}$. For our analysis, we set the barrier height and radius of the circular-box trap as $V_{0b}=100$ and $R_{\mathrm{box}}=6$, respectively. The crust potential, which acts as a vortex pinning site, is given by [15, 14]:
$V_{\mathrm{crust}}(\mathbf{r},t)=V_{0c}\exp\left(-\frac{(|\mathbf{r}_{p}|-r_{\mathrm{crust}})^{2}}{(\Delta r_{\mathrm{crust}})^{2}}\right)\tilde{V}(x_{\theta},y_{\theta}),$ (4)
where $\tilde{V}(x_{\theta},y_{\theta})=3+\cos(n_{\mathrm{crust}}x_{\theta})+\cos(n_{\mathrm{crust}}y_{\theta})$, and the rotated coordinates are $x_{\theta}=\cos(\theta(t))x_{p}+\sin(\theta(t))y_{p}$ and $y_{\theta}=-\sin(\theta(t))x_{p}+\cos(\theta(t))y_{p}$. Here, $n_{\mathrm{crust}}$ determines the number of pinning sites, $r_{\mathrm{crust}}$ is the radius at which $V_{\mathrm{crust}}$ is maximized, and $\Delta r_{\mathrm{crust}}$ represents the crust thickness. The crust potential parameters are based on Refs. [15, 14]. The centrifugal potential, $V_{\mathrm{cent}}(\mathbf{r},t)=\frac{1}{2}\Omega(t)^{2}r^{2}$, ensures uniform condensate density under rotation. Following Refs. [33, 36, 37], we adopt a coherent length scale (comparable to vortex core size) of $\xi=10~\mathrm{fm}$, yielding an energy scale of $\epsilon=207~\mathrm{keV}$. With these parameters, the time scale is calculated as $\tau=3.2\times 10^{-21}~\mathrm{s}$. We compute the kinetic energy components and their spectra based on the analysis of their numerical implementation in Ref. [38]. It is important to note that while the GPE model is inherently limited to describing weakly interacting BECs in relatively small systems, resulting in a mismatch with the microscopic length scales present in neutron stars, it effectively captures vortex depinning phenomena, which are considered strong candidates for characterizing glitch events. Although scaling the vortex population, condensate radius and rotation frequency to astrophysical scales is unfeasible, the GPE model has reproduced Self-Organized Criticality of the pulsar glitch behavior  [16, 15] The simulation involves numerically solving the GPE using the split-step Crank-Nicolson method [39, 40] in a computational domain of size $512\times 512$, with spatial steps $dx=dy=0.05$ and a time step $dt=10^{-3}$ to ensure numerical stability and convergence. The numerical simulations are performed on an NVIDIA A100 GPU using CUDA C codes developed based on [41]. Initially, we prepare the condensate and crust with an initial rotation frequency $\Omega_{0}=2$ through imaginary-time iterations ($t\to-\mathrm{i}t$). This converged profile is then spun down at varying rates, characterized by the spin-down time $t_{s}$, in real-time. We first consider a case with a static crust and no dissipation ($\gamma=0$) while exploring the effects of damping in subsequent analyses.
Results: Superfluidity significantly impacts the interior dynamics of neutron stars, forming quantized vortices under rotation. These vortices facilitate angular momentum transfer, driving glitch behavior, while the superfluid density provides a qualitative picture of quantized vortices during the spin-down process. The density profiles shown in Fig. 1 indicate that the centrifugal potential initially dominates over the circular box trap due to the high rotation frequency.
Figure 1: Snapshots of condensate density during real-time spin-down at $t_{s}=10$ with the crust potential. As the condensate spins down, the initially dominant centrifugal confinement [(a)] is overtaken by the circular box trap [(b)–(d)]. Concurrently, turbulent flow induces the depinning of vortices.
As the condensate spins down, the influence of the box trap becomes dominant, resulting in a more uniform density distribution. The presence of the crust significantly increases the vortex number by providing pinning sites within the condensate and also accelerates vortex decay through a depinning mechanism. Liu and colleagues [33, 10] attribute this behavior to the Magnus flow generated around the pinning sites of the crust.
The Magnus flow, induced by the spin-down of the condensate, reaches a critical velocity that triggers vortex depinning and simultaneous density fluctuations. As the condensate approaches the spin-down time $t_{s}$, the flow enters a turbulent regime driven by this critical velocity. After spin-down, turbulent fluctuations persist briefly before decaying due to the absence of external forcing.
Further, the vortex avalanche triggered by Magnus flow results in an instantaneous recoupling of the superfluid and normal components, accompanied by a rapid transfer of angular momentum. This abrupt transition in superfluid systems may induce turbulence but cannot be confirmed using density profiles alone. To characterize the turbulent flow in the system, we therefore analyze the incompressible kinetic energy spectra of the condensate.
Figure 2: Incompressible kinetic energy spectra exhibiting the Kolmogorov cascade for spin-down time  $t_{s}=10$ (a), averaged over $t=10$ to $t=30$; (b) averaged over $t=30$ to $t=50$, and spin-down time  $t_{s}=20$ (c) averaged over $t=20$ to $t=50$; and (d) averaged over $t=50$ to $t=60$. The spectra initially [(a) and (c)] and later durations [(b) and (d)] exhibit $k^{-5/3}$ and $k^{-1}$ scalings, respectively alongside a $k^{-3}$ scaling.
We characterize the turbulent regimes by using Thomas-Fermi radius $R_{TF}$, the inter-vortex distance $\ell_{0}$, and the healing length $\xi$. For two-dimensional systems, the inter-vortex distance is approximated as $\ell_{0}=1/\sqrt{l_{v}}$, where $l_{v}$ represents the vortex density per unit area.
In the turbulent regime, the incompressible kinetic energy spectrum exhibits a $k^{-5/3}$ scaling in the inertial range $2\pi/R_{TF}<k<2\pi/\ell_{0}$ [see Figs. 2(a) and 2(c)], associated with vortex breakdown. Additionally, the condensate enstrophy undergoes a self-similar cascade, manifested as a $k^{-3}$ power-law scaling at large $k$ values in the incompressible spectrum. For longer spin-down times, the $k^{-5/3}$ scaling weakens, while the $k^{-3}$ scaling becomes more pronounced. The observation that turbulent instability is triggered by rapid transitions in angular frequency and that the energy cascade extends beyond the intervortex spacing aligns with findings on rotational two-dimensional quantum turbulence under varying rotation frequencies [23].
Following the initial turbulent behavior exhibiting a Kolmogorov cascade, the vortex system transitions to a Vinen turbulence phase as several vortices decay due to the spin-down. This Vinen turbulence is characterized by a $k^{-1}$ power-law scaling in the range $k>2\pi/\ell_{0}$ [see Fig. 2(b) and 2(d)], where individual vortices dominate the velocity field. Notably, the intervortex spacing decreases as vortex density declines, enabling the $k^{-1}$ scaling to manifest at these length scales. Similar to the Kolmogorov regime, the $k^{-1}$ scaling is more pronounced for shorter spin-downs (smaller $t_{s}$ values).
As the condensate spins down, the vortex population decreases, reducing the vortex-line density $l_{v}$. For 2D systems, $l_{v}$ can be approximated by computing the vortex population over a given area, which leads to increased intervortex spacing $\ell_{0}$ from the Kolmogorov to the Vinen regime.
Figure 3: Variation of vortex density (per unit area) with respect to time on a logarithmic scale for spin-down times (a) $t_{s}=10$, (b) $t_{s}=20$, (c) $t_{s}=30$, and (d) $t_{s}=40$. The vortex decay in the Kolmogorov and Vinen turbulence regimes exhibits $t^{-3/2}$ and $t^{-1}$ scaling behaviors, respectively.
As shown in Fig. 3, the vortex density exhibits a distinct scaling behavior with respect to time, further confirming the turbulent behavior of the condensate. During the initial Kolmogorov turbulent regime, the vortex-line density $l_{v}$ scales as $t^{-3/2}$, transitioning to $t^{-1}$ scaling in the later Vinen turbulent regime. This temporal decay of vortex-line density is attributed to vortex breakdown in the Kolmogorov regime, as well as to sound radiation (Kelvin-wave emission) and vortex reconnection in the Vinen regime [42, 43, 44]. For shorter spin-down times, the $t^{-3/2}$ scaling persists over a longer timescale, corresponding to a broader spatial extent of the $k^{-3/2}$ scaling in the incompressible kinetic energy spectrum (Fig. 2). Due to the intrinsic scales of the Gross–Pitaevskii framework, the power-law behavior observed here is limited to laboratory-scale systems. However, the functional form of this scaling, the $k^{-5/3}$ spectrum, is consistent with the incompressible turbulence spectrum reported in a distinct astrophysical context [25, 26]. This similarity in spectral behavior, despite the vast difference in physical scales and systems, suggests a common underlying turbulent mechanism may be at play.
Despite the spatial and temporal profiles of energy and vortex density indicating turbulent behavior, the onset of turbulence after the condensate spins down requires further investigation. We analyze the temporal behavior of the kinetic energy components and observe that the incompressible kinetic energy (associated with vortex flow) initially dominates but decreases as the condensate spins down without rotational forcing to inject vortices into the system. The quantum pressure energy, which is not associated with the velocity flow of the condensate, becomes dominant over the incompressible component. The spin-down time marks the crossover point where the quantum pressure component surpasses the incompressible component, beyond which Kolmogorov turbulence is observed in the condensate. This finding contrasts with turbulent decay in self-gravitating condensates, as reported in Ref. [26], where the dominance of the quantum pressure component indicates the absence of turbulence. This discrepancy is attributed to differences in the turbulent decay mechanisms. In the self-gravitating case, the decay of incompressible kinetic energy results from the instantaneous expulsion of vortex structures, whereas in the spin-down case, it arises from more gradual vortex decay. Despite reaching the spin-down time, vortices persist in the condensate (see Fig. 1) for a sufficient duration to undergo an energy cascade and exhibit turbulent behavior.
Figure 4: Temporal profiles of the total kinetic energy exchange between its components for spin-down times(marked by dark-grey vertical lines) (a) $t_{s}=10$, (b) $t_{s}=20$, (c) $t_{s}=30$, and (d) $t_{s}=40$. The red dashed line represents energy transfer from the incompressible to the compressible component ($E_{\mathrm{ex}}^{ic}$), the green dash-dotted line represents energy transfer between the incompressible and quantum pressure components ($E_{\mathrm{ex}}^{iq}$), and the blue line with circles represents energy transfer between the compressible and quantum pressure components ($E_{\mathrm{ex}}^{cq}$). A negative exchange energy indicates a transfer from the first component $\alpha\in\{i,c,q\}$ in the superscript to the second $\beta\in\{i,c,q\}$, while a positive value indicates the reverse.
Although glitch events are explained by the transfer of angular momentum [45], analyzing the onset of turbulence and the dynamics of quantized vortices is required to fully understand the transport of kinetic energy components, which, in turn, enables a new perspective on vortex depinning. The evolution of net energy transport between the kinetic components, shown in Fig. 4, provides critical insight into the post-spin-down dynamics and explains the persistence of vortex flow. The energy exchange between the compressible–quantum pressure and incompressible–compressible components does not significantly influence the turbulent dynamics. However, the energy exchange between the incompressible and quantum pressure components, denoted $E_{\mathrm{ex}}^{iq}$, is initially negative, indicating energy transfer from the incompressible to the quantum pressure component. After the condensate spins down, the exchange becomes positive, indicating a transfer from quantum pressure to incompressible component and at later times $E_{\mathrm{ex}}^{iq}\to 0$, as no further incompressible energy is transferred to quantum pressure. For sufficiently short spin-down times [see Fig. 4(a)–(b)], the exchange energy becomes positive, indicating an injection of incompressible energy from the quantum pressure component. This secondary injection mechanism sustains turbulent behavior in the condensate, even after the loss of rotational forcing, which explains the existence of vortex flow after the spin-down time is reached. For longer spin-down times, the peak of the secondary injection diminishes, reducing the intensity of turbulence.
The damping coefficient $\gamma$ is used in atomic BECs to model vortex-sound interactions [46]. A comparable approach is employed in neutron star interiors to describe interactions with the outer crust and proton superconducting core. In addition, the impact of $\gamma$ on vortex and turbulent dynamics under the dGPE model provides an important clue to the contribution of the interaction between neutron star interiors, which are mainly composed of superfluid and proton superconductors. For atomic BECs, the damping parameter $\gamma$ can be controlled by varying the temperature of the condensate, significantly affecting the vortex and density-wave dynamics. The damping parameter $\gamma$ is controlled by varying the temperature of the condensate, significantly affecting vortex and density-wave dynamics.
Figure 5: Variation of (a) incompressible kinetic energy spectra and (b) compressible kinetic energy spectra for different values of the damping coefficient $\gamma$. Both spectra are calculated for the case $t_{s}=10$ and averaged over the time range $t=10$ to $t=30$. The spectra exhibit a better fit to turbulent scaling for smaller $\gamma$ values.
Figure 6: Time evolution of the exchange energy between the incompressible and quantum pressure components for spin-down time $t_{s}=10$ (dark-gray vertical line). The different curves represent the behavior of the profile under varying damping strengths $\gamma$. Notably, the injection of incompressible energy is maximized for an optimal damping coefficient, corresponding to neither underdamped nor overdamped conditions.
Figure 5 illustrates the evolution of incompressible and compressible kinetic energy spectra for various values of the damping coefficient $\gamma$. For the strongly turbulent case with $t_{s}=10$ [see Figs. 2(a)], the spectra exhibit pronounced $k^{-5/3}$ scaling in the incompressible component and $k$ scaling in the compressible component for smaller damping values (underdamped regime). The $k$ scaling in the infrared region and $k^{-7/2}$ scaling in the ultraviolet regime for the compressible spectra indicates the presence of weak wave turbulence [21, 24, 26]. At approximately $\gamma\approx 10^{-2}$, the spectra abruptly lose their scaling behavior as stronger damping (overdamped regime) dissipates the turbulent flow. This turbulent flow is further confirmed by the emergence of a prominent $k^{-3}$ scaling and the diminishing $k^{-5/3}$ scaling in the overdamped case ($\gamma\approx 10^{-2}$), indicating the inhibition of vortex breakdown and the formation of a stable vortex lattice structure. It is worthwhile to mention mainly an over-damped condensate ($\gamma>0.02$) has been considered to model superfluid cores of neutron stars due to several potential dissipation pathways [13, 9].
Similar to the non-damped case, we complement the spectral analysis with the exchange energy profile in the presence of damping Fig. 6. The injection of incompressible kinetic energy reaches a maximum at an optimal value of $\gamma$, situated between the underdamped and overdamped regimes. Under these optimal damping conditions, the transfer of energy from quantum pressure to incompressible modes appears to be significantly enhanced. Given that $\gamma$ serves as a simplified parametrization of dissipation in quantum fluids, the underlying mechanism responsible for the observed maximum in incompressible energy injection at this intermediate damping requires further investigation. In the overdamped regime, around $\gamma\approx 10^{-2}$, the positive peak in exchange energy is absent, consistent with the suppression of turbulence under such strongly dissipative conditions.
Summary and Conclusion: We investigated two-dimensional quantum turbulence in a spinning-down condensate subjected to a crust potential that provides vortex pinning sites. By analyzing a continuous spin-down profile, we demonstrated that as the spin-down duration $t_{s}$ approaches the quasi-discrete limit, the condensate exhibits turbulent behavior. Following the spin-down duration $t_{s}$, the neutron star core displays turbulence characterized by an incompressible kinetic energy spectrum with a $k^{-5/3}$ scaling, indicative of the Kolmogorov regime, and a compressible energy spectrum with a $k^{1}$ scaling, consistent with thermalization. The turbulent core subsequently transitions to the Vinen regime, marked by a $k^{-1}$ scaling, where isolated vortices dominate the flow. In both regimes, the scaling behavior weakens as the spin-down duration increases. The vortex-line density $l_{v}$ exhibits a power-law decay, scaling as $t^{-3/2}$ in the Kolmogorov regime and as $t^{-1}$ in the Vinen regime.
The interplay between kinetic energy components reveals that spinning down induces an energy transfer from the incompressible component to quantum pressure. After the spin-down, a reverse transfer from quantum pressure to incompressible energy occurs, which is sufficient to trigger a turbulent cascade after the cessation of rotational forcing. This injection of kinetic energy is indirectly observed in the vortex-line decay profile as an increase in vortex-line density around $t_{s}$, resulting from the transfer of energy back into the incompressible component. This behavior is attributed to the depinning of vortex lattice, leading to a vortex avalanche [47, 9, 48].
The variation of the Kolmogorov spectrum and compressible spectrum with the damping parameter $\gamma$ shows a sharp transition from non-turbulent to turbulent scaling behavior occurs near $\gamma\sim 10^{-2}$. We have also observed an optimal damping coefficient, where the transfer from quantum pressure to incompressible energy is maximized. Under longer spin-down durations, Liu et al. [33] have reported pulsar glitches triggered by vortex avalanches. As quantum turbulence has been shown to suppress collective excitations in condensates [27, 28], the parameters determining the onset of turbulence presented here may be relevant to glitch formation in the presence of a crust potential. While vortex depinning is commonly considered the mechanism behind glitches in crust-bound systems, self-gravitating BECs without pinning potentials can also exhibit vortex expulsion to the periphery when the total circulation exceeds a critical threshold $\xi_{c}$ [49, 26]. Such expulsions could similarly interact with an external crust and trigger glitch-like behavior in the absence of depinning.
The onset of turbulence in BECs with pinning potentials during spin-down provides crucial insight into the prevalent GPE models of neutron star cores [15, 14, 33]. The dependence of turbulence dynamics on spin-down profiles and dissipative interactions within our Gross-Pitaevskii analog allows us to map the model’s parameter space. This mapping is crucial for assessing the theoretical feasibility and inherent limitations of the analogy itself, thereby clarifying which aspects of neutron star glitch phenomenology it may, or may not, be suited to explore [33]. Furthermore, this study is valuable for exploring inhomogeneous two-dimensional quantum turbulence and its associated dynamics, particularly in relation to the emergence of secondary energy injection mechanisms. It is worth noting that in systems with a large number of vortices, secondary energy injection is primarily governed by the transfer of energy from compressible to incompressible modes. This process arises due to vortices traveling longer distances, repeatedly pinning and depinning at various locations. Similar to the partial spin-down scenario, this results in a substantial retention of vortices within the condensate, which inhibits the transfer of energy from quantum pressure to incompressible kinetic modes.
In conclusion, although the Gross-Pitaevskii analog necessarily simplifies the fermionic nature of neutron Cooper-pair superfluidity in neutron-star interiors—including density-dependent 1S0 and 3P2 pairing, possible hyperon contributions, and uncertainties in nuclear many-body physics—it remains the most microscopically faithful effective theory for studying the dynamics of individual quantized vortices and their mutual interactions. While directly applicable only to finite, laboratory-scale systems, it provides essential insight into the fundamental quantum mechanisms that may underpin vortex-mediated phenomena in astrophysical superfluids [33]. These additional physical contributions could, in principle, be incorporated through multi-component or binary-fluid extensions of the GP framework, which offers a promising direction for future research.
Our simulations reveal a previously unreported transient anomalous energy-injection mechanism, driven by quantum pressure, that operates during the highly turbulent phase immediately following major vortex avalanches. Although quantitative extrapolation to kilometre-scale neutron stars is not yet possible, and the effect diminishes under prolonged differential rotation, the consistent appearance of avalanche-like events, power-law glitch-size distributions, and sub-second rise times across numerous independent Gross-Pitaevskii studies supports the astrophysical relevance of the underlying vortex dynamics.
The anomalous energy-injection mechanism identified in the Gross–Pitaevskii analog system, therefore, represents a new and potentially relevant ingredient for future multi-scale and hybrid models of neutron star interiors.
Acknowledgements.
A.S. acknowledges financial support from the Council of Scientific and Industrial Research (CSIR), India, in the form of a Direct Senior Research Fellowship. The work of P.M. is supported by the Ministry of Education-Rashtriya Uchchatar Shiksha Abhiyan (MoE RUSA 2.0): Bharathidasan University – Physical Sciences.